OpenStudy (anonymous):

3(n+1)! / 5n! Please show your steps!

OpenStudy (lgbasallote):

hint: \[(n+1)! \implies (n+1) \times n!\] does that help?

jimthompson5910 (jim_thompson5910):

Hint: n! = n*(n-1)(n-2)...3*2*1 so (n+1)! = (n+1)*n*(n-1)(n-2)...3*2*1

jimthompson5910 (jim_thompson5910):

and you can rewrite the remaining bit n*(n-1)(n-2)...3*2*1 as n! like lgbasallote did to get (n+1)! = (n+1)*n!

OpenStudy (anonymous):

the n in n! is not the same as the variable 'n' which is alone in this case?

jimthompson5910 (jim_thompson5910):

not sure what you mean

OpenStudy (anonymous):

can I cancel out the n!

jimthompson5910 (jim_thompson5910):

yes